# A unified duality framework for barotropic, quantum and Korteweg fluids

### v1

@@ 1. Introduction
In this paper, we study the variational dual formulation of several compressible fluid models. The approach originates with Brenier [9, 10], who proposed, for the incompressible Euler system, the compressible barotropic Euler system, the Burgers equation, and related models, to consider solutions minimizing the Lagrangian action. While this minimization may not always admit a solution, it naturally leads to a dual problem with more favorable convexity properties. His results imply the possibility of recovering smooth solutions of the compressible barotropic Euler system over small time intervals from the dual formulation, see Remark 5.2 below. This method was applied to the multi-stream Euler–Poisson system in [11], and a numerical implementation for the quadratic porous medium and Burgers equations was recently proposed in [55]. Further developments of Brenier’s approach, presented in [61, 62], reveal structures in many nonlinear PDEs that allow variational dual formulations with advantageous properties. These dual problems are closely connected to optimal transport, specifically its ballistic variant [61, 62, 55]. A similar variational dual formulation has been proposed in [1, 2, 58] and related works. A central feature of the latter framework is the use of a base state, which serves as an “initial guess” for the PDE solution; the entropy is then replaced by an appropriate relative entropy. In this view, the frameworks of [9, 10, 61, 62, 55] use zero base states. The recent preprints [3, 63] build upon both approaches; among other contributions, [63] introduces a duality-based gradient flow scheme for approximating solutions of PDEs for which the existence of global weak solutions may not yet have been established. In [62], we extended Brenier’s duality formalism to systems of PDEs that can be written in the abstract form
@@
In Section 2, we introduce a general abstract framework that unifies all three systems, as well as numerous other models from [9, 61, 62], enabling their simultaneous analysis. In Section 3, by incorporating time-dependent weights, we establish the consistency of the duality scheme over large time intervals (Theorem 3.2 and Remark 3.3). As an application, we formulate and prove a version of the “Dafermos principle” (Theorem 3.5) for these models: no subsolution can dissipate the total entropy sooner or at a faster rate than a suitably defined strong solution within the lifespan of the latter. In Section 4, we demonstrate the solvability of the corresponding dual problems, for continuous vacuum-free initial data, in spaces of finite Radon measures, and show that no duality gap arises (Theorem 4.1). Section 5 provides a detailed explanation of how the compressible fluid systems mentioned above fit into our framework. In Section 6, we revisit Brenier’s treatment [9] of the duality associated with the inviscid Burgers equation and clarify certain points that may have remained ambiguous in his original work. While Appendix A is purely technical, Appendix B concludes the paper by highlighting several challenging open questions.
@@ 2. Basic framework
In specific applications, reasonable “strong” solutions may exhibit a different regularity. Here we aim to work with a notion of strong solution that is as weak as possible while still ensuring the properties needed below. To fix the ideas, we restrict this definition to functions taking values in $\mathcal{O}$. In contrast, our (continuous) weak solutions are allowed to take values outside $\mathcal{O}$. We expect that the stronger constraint imposed on strong solutions is not essential and can be removed or relaxed with additional technical effort, including appropriate modifications of the definitions.
@@ 4. Existence and absence of a duality gap
In this section, we establish [^2] See, however, Open Problem B.1. the existence of solutions to the dual problem, which may be called variational dual solutions to (2.2)–(2.4) (cf. [58, 3, 63]), and demonstrate that no duality gap arises between $\tilde{\mathcal{I}}(v_{0},T)$ and $\tilde{\mathcal{J}}(v_{0},T)$. Note that here we obtain only a restricted form of absence of a duality gap: unlike Theorem 3.2, we cannot claim equality between $\mathcal{I}(v_{0},T)$ and $\tilde{\mathcal{I}}(v_{0},T)$.

### v2

@@ 1. Introduction
In this paper, we study the variational dual formulation of several compressible fluid models. The approach originates with Brenier [9, 10], who proposed, for the incompressible Euler system, the compressible barotropic Euler system, the Burgers equation, and related models, to consider solutions minimizing the time-integrated entropy. While this minimization may not always admit a solution, it naturally leads to a dual problem with more favorable convexity properties. His results imply the possibility of recovering smooth solutions of the compressible barotropic Euler system over small time intervals from the dual formulation, see Remark 5.2 below. This method was applied to the multi-stream Euler–Poisson system in [11], and a numerical implementation for the quadratic porous medium and Burgers equations was recently proposed in [55]. Further developments of Brenier’s approach, presented in [61, 62], reveal structures in many nonlinear PDEs that allow variational dual formulations with advantageous properties. These dual problems are closely connected to optimal transport, specifically its ballistic variant [61, 62, 55]. A similar variational dual formulation has been proposed in [1, 2, 58] and related works. A central feature of the latter framework is the use of a base state, which serves as an “initial guess” for the PDE solution; the entropy is then replaced by an appropriate relative entropy. In this view, the frameworks of [9, 10, 61, 62, 55] use zero base states. The recent preprints [3, 63] build upon both approaches; among other contributions, [63] introduces a duality-based gradient flow scheme for approximating solutions of PDEs for which the existence of global weak solutions may not yet have been established. In [62], we extended Brenier’s duality formalism to systems of PDEs that can be written in the abstract form
@@
In Section 2, we introduce a general abstract framework that unifies all three systems, as well as numerous other models from [9, 61, 62], enabling their simultaneous analysis. In Section 3, by incorporating time-dependent weights, we establish the consistency of the duality scheme over large time intervals (Theorem 3.2 and Remark 3.3). As an application, we formulate and prove a version of the “Dafermos principle” (Theorem 3.5) for these models: no subsolution can dissipate the total entropy sooner or at a faster rate than a suitably defined strong solution within the lifespan of the latter. In Section 4, we demonstrate the solvability of the corresponding dual problems, for continuous vacuum-free initial data, in spaces of finite Radon measures, and show that no duality gap arises (Theorem 4.1). Section 5 provides a detailed explanation of how the compressible fluid systems mentioned above fit into our framework. In Section 6.2, we revisit Brenier’s treatment [9] of the duality associated with the inviscid Burgers equation and clarify certain points that may have remained ambiguous in his original work. While Appendix A is purely technical, Appendix B concludes the paper by highlighting several challenging open questions.
@@ 2. Basic framework
It is obvious that a smooth solution to (2.2)–(2.4) with values in $\mathcal{O}$ is a strong solution in the sense of Definition 2.14. We aim to work with a notion of strong solution that is as weak as possible while still ensuring the properties needed below. To fix the ideas, we restrict this definition to functions taking values in $\mathcal{O}$. In contrast, our (continuous) weak solutions are allowed to take values outside $\mathcal{O}$. We expect that the stronger constraint imposed on strong solutions is not essential and can be removed or relaxed with additional technical effort, including appropriate modifications of the definitions.
@@ 4. Existence and absence of a duality gap
In this section, we establish [^2] See, however, Open Problem B.1. the existence of solutions to the dual problem, which may be called variational dual solutions to (2.2)–(2.4) (cf. [58, 3, 63]), and demonstrate that no duality gap arises between $\tilde{\mathcal{I}}(v_{0},T)$ and $\tilde{\mathcal{J}}(v_{0},T)$. Note that here we obtain only a restricted form of absence of a duality gap: unlike Theorem 3.2 and Remark 6.2, we cannot claim equality between $\mathcal{I}(v_{0},T)$ and $\tilde{\mathcal{I}}(v_{0},T)$.
@@ 6. Inviscid Burgers’ equation: revisiting Brenier’s shock-free substitute
Remark 6.2.
Let us verify that, in the context of this section, there is no gap between $\mathcal{I}(v_{0},T)$ and $\tilde{\mathcal{J}}(v_{0},T)$. By (4.2), $\tilde{\mathcal{I}}(v_{0},T)=\tilde{\mathcal{J}}(v_{0},T)$, and it follows from [9, Proposition 4.4] that $\tilde{\tilde{\mathcal{I}}}(v_{0},T)=\tilde{\mathcal{J}}(v_{0},T)$, where $\tilde{\tilde{\mathcal{I}}}$ is defined as in (3.3), but restricted to smooth pairs $(v,M)$. Thus it is enough to show that $\mathcal{I}(v_{0},T)\leq\tilde{\tilde{\mathcal{I}}}(v_{0},T)$. This follows from the observation that the time integral of the total entropy $K(t)$ of the entropy solution $v$ cannot be larger than the time integral of the total entropy $\tilde{K}(t)$ of any smooth subsolution $(u,M)$ to (6.1). Indeed, for every such subsolution there exists a smooth function $\theta:[0,T]\times\mathbb{T}^{1}\to\mathbb{R}$ such that
$\partial_{x}\theta=u,\ \partial_{t}\theta+\frac{1}{2}M=0,\ \theta(0,\cdot)=\varphi,\ M\geq u^{2}.$
By comparison principle, $\theta\leq\psi$, where $\psi$ is the viscosity solution to the quadratic Hamilton-Jacobi equation emanating from $\varphi$. Consequently,
$\int_{0}^{T}K(t)\,dt=\frac{1}{2}\int_{0}^{T}(v,v)\,dt=\frac{1}{2}\int_{0}^{T}(\partial_{x}\psi,\partial_{x}\psi)\,dt\int_{\mathbb{T}^{1}}=\int_{\mathbb{T}^{1}}\left[\varphi(x)-\psi(T,x)\right]\,dx\\ \leq\int_{\mathbb{T}^{1}}\left[\varphi(x)-\theta(T,x)\right]\,dx=\int_{0}^{T}\tilde{K}(t)\,dt.$
@@ Appendix B Some open problems
Acknowledgement
The work was partially supported by CMUC [^4] https://doi.org/10.54499/UID/00324/2025 under FCT, grants UID/00324/2025 and UID/PRR/00324/2025, and by KAUST Research Funding (KRF), award ORFS-2024-CRG12-6430.3. The author thanks L.-P. Chaintron for helpful discussions.